The direct materials and direct labour standards for one bottle of Clean-All spray cleaner are given below:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
  Direct materials 5.0   millilitres $ 0.30 per millilitre $ 1.50  
  Direct labour 0.4   hours $ 12.00 per hour $ 4.80  

During the most recent month, the following activity was recorded:

  1. 28,000 millilitres of material was purchased at a cost of $0.25 per millilitre.
  2. All of the material was used to produce 3,000 bottles of Clean-All.
  3. 675 hours of direct labour time was recorded at a total labour cost of $8,100.

Required:

1. Compute the direct materials price and quantity variances for the month. (Indicate the effect of each variance by selecting "F" for favourable, "U" for unfavourable, and "None" for no effect (i.e., zero variance).)

2. Compute the direct labour rate and efficiency variances for the month. (Indicate the effect of each variance by selecting "F" for favourable, "U" for unfavourable, and "None" for no effect (i.e., zero variance).)
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (0.30-0.25)*28,000

= 1400 F

Materials Quantity Variance = (SQ-AQ) *SP

= (3000*5 - 28,000)*0.30

= 3900 U

Labor Rate Variance = (SR-AR) *AH

= (12 - 8100/675) * 675

= 0

Labor Efficiency variance = (SH-AH) *SR

= (3000*0.4 - 675)*12

= 6300F

Comment if you face any issues